### v1

@@ Abstract
Synthetic data sets are used in cosmology to test analysis procedures, to verify that systematic errors are well understood and to demonstrate that measurements are unbiased. In this work we describe the methods used to generate synthetic datasets of Lyman-$\alpha$ quasar spectra aimed for studies with the Dark Energy Spectroscopic Instrument (DESI). In particular, we focus on demonstrating that our simulations reproduces important features of real samples, making them suitable to test the analysis methods to be used in DESI and to place limits on systematic effects on measurements of Baryon Acoustic Oscillations (BAO). We present a set of mocks that reproduce the statistical properties of the DESI early data set with good agreement. Additionally, we use full survey synthetic data to forecast the BAO scale constraining power with DESI.
@@ 1 Introduction
The Lyman-$\alpha$ forest is a series of absorption features present in quasar [^1] Also referred indistinctly as quasi-stellar object or QSO. spectra caused by intervening neutral hydrogen (HI) clouds along the line of sight and has been proven to be a powerful cosmological tool due to the tight relation between the Lyman-$\alpha$ optical depth and the densities of gas and dark matter. Lyman-$\alpha$ absorption observed in spectra of distant quasars is now widely used to: (i) calculate 3D correlation functions to study Baryon Acoustic Oscillations (BAO) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]; (ii) measure the line-of-sight (1D) power spectrum to constrain the amplitude and shape of the matter power spectrum at high redshift [11, 12, 13], and thus to yield tight constraints on neutrino masses [14, 15, 16] and dark matter models [16, 17, 18, 19, 20]; (iii) study the evolution of the intergalactic medium (IGM) through techniques such as tomography [21, 22].
@@
For all of these studies, as in general cosmological measurements, it is necessary to characterize possible sources of systematic effects as well as to test analysis pipelines. In this regard, the use of realistic synthetic data sets has acquired an important role over the years. Lyman-$\alpha$ synthetic spectra for BAO analysis and methods to generate them have been used since the analysis of the Baryon Oscillation Spectroscopic Survey (BOSS) year 1 data [1] and later on in BOSS Data Release 9 (DR9) [2, 3], which used mock data sets following the prescription in [23]. Later, the analysis of BOSS DR11 [5] and DR12 [6] used improved mock data developed by [24] which included astrophysical effects such as absorption due to transition lines different from Lyman-$\alpha$. These mocks also used the method of [25] to produce a realistic Lyman-$\alpha$ – QSO cross-correlation and were used to validate the analysis of [7]. The most recent effort on producing Lyman-$\alpha$ mocks was during the analysis presented in extended Baryon Oscillation Spectroscopic Survey (eBOSS) DR16 [10] including the use of two sets of simulations: LyaCoLoRe and Saclay [^2] The name makes reference to the institution where these mocks were mainly developed. mocks. Simulated maps of Lyman-$\alpha$ flux transmission were produced by these two independent methods as described in [26] and [27] respectively. They were both post-processed, using the methodology described in this work, to produce the final synthetic spectra. In all these cases, the role of synthetic quasar spectra has been to test the analysis methodology, to study systematics, and ultimately to validate the BAO measurements.
@@
The currently running Dark Energy Spectroscopic Instrument (DESI) is expected to measure the spectra of 40 million galaxies and over a million Lyman-$\alpha$ quasars with redshift $z>2$ in a 14,000 $\rm deg^{2}$ area during a five-year period, greatly improving the precision of the BAO scale measurements below an accuracy of 1% [28, 29]. Achieving this goal requires a robust characterization of systematic errors, necessitating the production of realistic Lyman-$\alpha$ synthetic spectra that accurately capture the statistical properties of the observed data, as well as the effects of astrophysical processes and instrumental noise.
@@
This manuscript is organized as follows: section 2 describes the overall methods used to produce synthetic Lyman-$\alpha$ spectra by using the transmitted flux fraction, quasar unabsorbed spectrum (continuum) template generation and the addition of astrophysical effects, while section 3 explains the strategies we follow to effectively reproduce a survey in terms of footprint, object density, redshift-magnitude distributions and instrument response model. In section 4 we present a set of 40 synthetic Lyman-$\alpha$ spectra catalogs that emulate the DESI Early Data Release (EDR) [30] plus 2 months of observations (DESI-M2) and present a qualitative comparison with observed data. In section 5 we take advantage of the results of our EDR+M2 mocks to study the use of mocks as a forecast tool for the DESI experiment and predict the ultimate BAO scale constraining power of DESI using full DESI mocks.
@@ 2 Simulating Lyman-$\alpha$ spectra
Simulating Lyman-$\alpha$ spectra with quickquasars is performed in three main stages. (i) The transmitted flux fraction is read from an input mock Lyman-$\alpha$ forest dataset, hereafter referred as raw mocks, e.g. the output of the LyaCoLoRe [26] or Saclay [27] programs. Optionally, the transmitted flux fraction can be combined with absorptions from Damped Lyman-$\alpha$ Absorbers (DLAs), Broad Absorption Lines (BALs), and additional absorbers, hereafter referred as metals. (ii) A template of the quasar unabsorbed spectrum is defined and multiplied by the transmitted flux fraction to create the noiseless quasar template. (iii) The instrumental model and observing conditions are introduced to simulate the noisy spectra. The first two of these stages are described in detail in the following sections, while the last will be described in section 3.
@@ 2.1 Transmitted flux fraction and raw mocks
These mocks were used for the Lyman-$\alpha$ BAO studies in eBOSS DR16 [10] and were referred as the London mocks. The methodology to produce this set of raw mocks is fully described in [26]. It starts with the generation of a correlated Gaussian multivariate overdensity field sampled in a three-dimensional mesh using the Cosmological Lofty Realizations package (CoLoRe [^6] https://github.com/damonge/CoLoRe[35], given an input matter power spectrum. The positions of the quasars in the mocks are assigned following an input number density $n(z)$ and bias $b(z)$, and placed in density peaks of the Gaussian random field above a given threshold via Poisson sampling. Density fields along the line of sight from the quasars to the center of the box, namely the skewers, are also extracted at this stage. The results are then post-processed with the LyaCoLoRe [^7] https://github.com/igmhub/LyaCoLoRe code, where small-scale fluctuations are added to the skewers in order to reproduce the variance of the Lyman-$\alpha$ forest in the data. A log-normal transformation is applied to the final Gaussian density field and the Fluctuating Gunn-Peterson approximation (FGPA) [36] is used to compute the optical depth on each simulated cell. Then, the effects of RSD are added by computing the radial velocity from the gradient of the Newtonian gravitational potential of the density field, and this velocity is then used to shift the position of the optical depth in redshift space accordingly. Lastly, the optical depth $\tau$ is transformed into the transmitted flux with the equation $F=e^{-\tau}$. In addition, the LyaCoLoRe post-processing method includes the generation of transmission due to metals and also provides a catalog with the redshift and column densities of DLAs that are correlated with the density field, as briefly described in section 2.3.
@@ 2.2 Continuum templates
  • SIMQSO (default): Templates are generated using the simqso library [^9] https://github.com/imcgreer/simqso [40, 41], which contains a very broad set of tools to generate mock quasar spectra. The model used has two main components: a broken power-law continuum, and a set of Gaussian emission lines defined by their wavelength, equivalent width, and Gaussian RMS width ($\sigma$). To emulate the power-law continuum component, we randomly generate the slopes of the broken power-law continuum model following a Gaussian distribution centered at the slope value $m$ with dispersion $\sigma_{m}$ in a particular rest-frame wavelength region. We use the default values from simqso outside the Lyman-$\alpha$ region which are based in BOSS DR9: $m=-0.37$ at $5700\ \text{\AA}<\lambda<9730\ \text{\AA}$, $m=-1.70$ at $9730\ \text{\AA}<\lambda<22300\ \text{\AA}$ and -1.03 at $\lambda>22300\ \text{\AA}$, all with dispersion $\sigma_{m}=0.3$. For the Lyman-$\alpha$ forest region we tuned the slopes and dispersion to better reproduce the mean continuum and the dispersion, from a PCA analysis, observed in eBOSS DR16 data [42]. The resultant slopes are: $m=-1.5$ at $\lambda<1100\ \text{\AA}$ and $m=-0.5$ at $1100\ \text{\AA}<\lambda<5700\ \text{\AA}$, both with dispersion $\sigma_{m}=0.7$. To emulate the emission line component, we use the model described in detail in appendix A. The model combines emission lines within the Lyman-$\alpha$ forest region from the composite model of BOSS spectra [43] and the emission lines outside the Lyman-$\alpha$ region from the simqso model. Emission line diversity is given mainly by the scatter in the equivalent width. Furthermore, the specific values for the mean equivalent width of some of the emission lines were adjusted so that the mean continuum resembles the results from DESI EDR+M2 data, as further discussed in section 4.2. Note that the emission line parameters used in this paper slightly differ from those used to generate the eBOSS DR16 mocks [10]. Information about the continuum template generated for each mock quasar, such as the slopes and emission lines equivalent widths, are stored in one of the output "truth" files from quickquasars which are described in appendix C.
@@ 2.3 Astrophysical contaminants
Observed quasar spectra include features due to different astrophysical effects that contaminate the correlation function. In order to study their impact these features may be added to the generated synthetic spectra. Below we describe these contaminants.
Broad Absorption Lines systems
Quasars with broad ultraviolet (UV) absorption in their spectrum are called Broad Absorption Line (BAL) quasars. This absorption is produced by surrounding gas clouds near the quasar nucleus. BAL quasars are further classified on the basis of spectral lines that show troughs typically blueshifted from their rest-frame wavelength and with velocity widths larger than $2000\,\rm{km/s}$ [46]. For instance, HiBALs are BALs with absorption in high-ionization lines such as C IV(1549), LoBALs are those with absorption troughs from low-ionization features such as Mg II, whereas FeLoBALs exhibit absorption in Fe lines such as Fe II [47, 48].
@@
High Columns Density systems
@@
  • Random. The simplest method is to add a random number of HCDs to the transmission at random wavelengths. In this case, the generated HCDs would not be not correlated with the density field that was used to produce the raw mocks, which makes this method unsuitable for our purposes and therefore is not used for any of the mocks produced for this work. However, this method is being used in Ohio mocks to study the effect of DLAs in P1D analyses.
@@
Contamination from additional absorption transitions
Besides neutral-hydrogen, the IGM contains other absorbers, which we refer to as metals. Because our data analysis protocol transforms wavelengths to redshifts assuming Lyman-$\alpha$ absorption, hydrogen and a metal at the same physical position give absorption at different apparent redshifts corresponding to an apparent distance separation of
@@
where $D_{H}$ is the Hubble distance and $\Delta\lambda=\lambda_{m}-\lambda_{Ly\alpha}$ is the wavelength separation of the metal transition ($\lambda_{m}$) with respect to Lyman-$\alpha$. The auto and cross-correlation functions will then show peaks at $r_{\perp}\approx 0$ and the $r_{\parallel}$ for each metal of sufficiently strong absorption. The most important such metals are Si II(1260), Si III(1207), Si II(1193), and Si II(1190) corresponding respectively to $r_{\parallel}=104,\,-21,\,-54$ and $-61{\ \mathrm{Mpc/h}}$ at $z=2.35$.
In addition to this effect due to metal-neutral-hydrogen correlations, particularly strong absorbers can significantly affect the measured auto-correlation through their own (metal-metal) auto-correlation. The most important transitions are C IV(1548) and C IV(1550). However, these are not produced in our mocks because the effect on the Lyman-$\alpha$ forest comes through C IV at redshifts as low as $z\approx 1.32$, outside the range of the transmission files of the raw mocks.
@@
  • Added by quickquasars. In this case we take the Lyman-$\alpha$ transmitted flux fraction at rest-frame as it is given in the transmission files. The optical depth of the Lyman-$\alpha$ absorption is then re-scaled by a factor $C_{m}$ according to the relative absorption strength of each metal with respect to Lyman-$\alpha$. The transmitted flux fraction of the metals is then computed by $F_{m}=e^{-C_{m}\tau_{Ly\alpha}}$. Finally, the absorptions are moved to their corresponding observed wavelengths by using the equation $\lambda_{obs,m}=(1+z_{abs})\lambda_{m}$, where $z_{abs}=\lambda/\lambda_{Ly\alpha}-1$ is the redshift at which the absorption occurs. In this method, the re-scaling of the optical depth is done after RSD effects have been applied to the Lyman-$\alpha$ absorptions forcing the metals to have the same RSD parameter as Lyman-$\alpha$.
    To find the value of $C_{m}$ we made several realizations of mocks including Lyman-$\alpha$ absorption and the contamination due to the transitions of Si II(1260), Si III(1207), Si II(1193), and Si II(1190), varying the value of $C_{m}$ in each realization. The absorption amplitude of each metal transition strongly affects the value for its corresponding transmitted flux fraction bias $b_{m}$ ($b_{i}$ defined in equation 2.1, for metals). In other words, the bias highly depends on the intensity of the optical depth of each independent transition and consequently on the value $C_{m}$ used in this method. For this simple simulation technique, $b_{m}$ and $C_{m}$ have a linear relation. This allowed us to find the $C_{m}$ needed to reproduce the measured biases of different eBOSS datasets. We chose the absorption coefficient value that reproduced the results from eBOSS DR14 [8]. This procedure was made using the Lyman-$\alpha$ transmission files from both, LyaCoLoRe mocks and Saclay mocks, the $C_{m}$ values found are shown in table 1, which are set as default in quickquasars when using this method. These values were used to compute the bias for a sample of ten LyaCoLoRe and ten Saclay mock realizations and were found to be consistent with the target bias value of each metal in DR14. We refer the reader to [57] and [58] for further details about these results. The same values were also used to produce the mocks presented in eBOSS DR16 analysis [10], also presenting consistency between mocks and observed data.
@@
    Si II(1190)
    $6.4239\times 10^{-4}$ $4.4960\times 10^{-4}$
    Si II(1193)
    $9.0776\times 10^{-4}$ $6.3540\times 10^{-4}$
    Si II(1260)
    $3.5420\times 10^{-4}$ $4.2504\times 10^{-4}$
    Si III(1207)
    $1.8919\times 10^{-3}$ $9.4595\times 10^{-4}$
  • From the transmission files. In this method the transmitted flux fraction of metals is computed during the generation of the Lyman-$\alpha$ raw mocks. This was done for LyaCoLoRe raw mocks using the methodology explained in section 5.2 of [26], which we briefly describe here. Similarly to the method followed in quickquasars, here we assume that the optical depth of each additional metal is proportional to that of the Lyman-$\alpha$ absorption. Therefore for each metal we need to define a relative absorption strength $A_{m}$, so that $\tau_{m}=A_{m}\tau_{Ly\alpha}$ and a rest-frame wavelength at which the metal absorption will be included. Notice that in general the absorption strength $A_{m}$ here is not the same as the coefficient $C_{m}$ mentioned above. However, the LyaCoLoRe mocks produced in [26] made use of the coefficients found using the quickquasars method, described above. [^11] All the relative absorption strength values used to produce the metal transmissions correspond to the $C_{m}$ coefficients found with the quickquasars method except for Si II(1190) where a value $A_{\text{Si~{}II(1190)}}=-9.0776\times 10^{-4}$ was used (see Table 2 in [26]). Finally, RSDs are applied and the transmitted flux fraction of each metal is computed separately. The aforementioned Si and Lyman-$\beta$ transmissions are then saved independently in the transmission files, which can then be read by quickquasars in the same way as the Lyman-$\alpha$ transmissions and added to the total transmitted flux.

### v3

@@ Abstract
Synthetic data sets are used in cosmology to test analysis procedures, to verify that systematic errors are well understood and to demonstrate that measurements are unbiased. In this work we describe the methods used to generate synthetic datasets of Lyman-$\alpha$ quasar spectra aimed for studies with the Dark Energy Spectroscopic Instrument (DESI). In particular, we focus on demonstrating that our simulations reproduces important features of real samples, making them suitable to test the analysis methods to be used in DESI and to place limits on systematic effects on measurements of Baryon Acoustic Oscillations (BAO). We present a set of mocks that reproduce the statistical properties of the DESI early data set with good agreement. Additionally, we use a synthetic dataset to forecast the BAO scale constraining power of the completed DESI survey through the Lyman-$\alpha$ forest.
@@ 1 Introduction
The Lyman-$\alpha$ forest is a series of absorption features present in quasar [^1] Also referred indistinctly as quasi-stellar object or QSO. spectra caused by intervening neutral hydrogen (HI) clouds along its line of sight and has been proven to be a powerful cosmological tool due to the tight relation between the Lyman-$\alpha$ optical depth and the densities of gas and dark matter. Lyman-$\alpha$ absorption observed in spectra of distant quasars is now widely used to: (i) calculate 3D correlation functions to study Baryon Acoustic Oscillations (BAO) [e.g. 1, 2, 3, 4, 5, 6, 7, 8, 9, 10]; (ii) measure the line-of-sight one-dimensional flux power spectrum ($P_{\rm{1D}}$) to constrain the amplitude and shape of the matter power spectrum at high redshifts [e.g. 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], and thus to yield tight constraints on neutrino masses [e.g. 21, 22, 23] and dark matter models [e.g. 24, 25, 26, 27, 28, 23], to mention some applications; (iii) study the evolution of the intergalactic medium (IGM) through techniques such as the tomography of the Lyman-$\alpha$ forest [e.g. 29, 30, 31, 32, 33, 34, 35, 36].
@@
For all of these studies, as in general cosmological measurements, it is necessary to characterize possible sources of systematic effects as well as to test analysis pipelines. In this regard, the use of realistic synthetic data sets has acquired an important role over the years. Lyman-$\alpha$ synthetic spectra for BAO analysis and methods to generate them have been used since the analysis of the Baryon Oscillation Spectroscopic Survey (BOSS) year one data [1] and later on in BOSS Data Release 9 (DR9) [2, 3], which used mock data sets following the prescription in [37]. Later, the analysis of BOSS DR11 [5] and DR12 [6] used improved mock data developed by [38] which included astrophysical effects such as absorption due to transition lines different from Lyman-$\alpha$. These mocks also used the method of [39] to produce a realistic Lyman-$\alpha$ – QSO cross-correlation and were used to validate the analysis of [7]. The most recent effort on producing Lyman-$\alpha$ mocks was during the analysis presented in extended Baryon Oscillation Spectroscopic Survey (eBOSS) DR16 [10] including the use of two sets of simulations: LyaCoLoRe and Saclay [^2] The name makes reference to the institution where these mocks were mainly developed. mocks. Simulated maps of Lyman-$\alpha$ flux transmission were produced by these two independent methods as described in [40] and [41] respectively. They were both post-processed, using the methodology described in this work, to produce the final synthetic spectra. In all these cases, the role of synthetic quasar spectra has been to test the analysis methodology, to study systematics, and ultimately to validate the BAO measurements.
@@
The currently ongoing Dark Energy Spectroscopic Instrument (DESI) [42, 43] survey is expected to measure the spectra of 40 million galaxies and over a million Lyman-$\alpha$ quasars with redshift $z>2$ in a 14,000 $\rm deg^{2}$ area during a five-year period, greatly improving the precision of the BAO scale measurements below an accuracy of 1% [42, 44]. Achieving this goal requires a robust characterization of systematic errors, necessitating the production of realistic Lyman-$\alpha$ synthetic spectra that accurately capture the statistical properties of the observed data, as well as the effects of astrophysical processes and instrumental noise.
@@
This manuscript is organized as follows: section 2 describes the overall methods used to produce synthetic Lyman-$\alpha$ spectra by using the transmitted flux fraction, quasar unabsorbed spectrum (continuum) template generation and the addition of astrophysical effects. Section 3 explains the strategies we follow to effectively reproduce a survey in terms of footprint, object density, redshift-magnitude distributions and instrument response model. Section 4 presents a set of 40 synthetic Lyman-$\alpha$ spectra catalogs that emulate the DESI Early Data Release (EDR) [45] plus 2 months of observations (DESI-M2) and present a qualitative comparison with observed data. Section 5 takes advantage of the results of our EDR+M2 mocks to study the use of mocks as a forecast tool for the DESI experiment and predict the ultimate BAO scale constraining power of DESI using mocks of the completed DESI Lyman-$\alpha$ survey sample.
@@ 2 Simulating Lyman-$\alpha$ spectra
Simulating Lyman-$\alpha$ spectra with quickquasars is performed in three main stages. (i) The transmitted flux fraction is read from an input mock Lyman-$\alpha$ forest dataset, hereafter referred as raw mocks, e.g. the output of the LyaCoLoRe [40] or Saclay [41] programs. Optionally, the transmitted flux fraction can be combined with absorption from Damped Lyman-$\alpha$ Absorbers (DLAs), Broad Absorption Lines (BALs), and additional absorbers, hereafter referred as metals. (ii) A template of the quasar unabsorbed spectrum is defined and multiplied by the transmitted flux fraction to create the noiseless quasar template. (iii) The instrumental model and observing conditions are introduced to simulate the noisy spectra. The first two of these stages are described in detail in the following sections, while the last will be described in section 3.
@@ 2.1 Transmitted flux fraction and raw mocks
These mocks were used for the Lyman-$\alpha$ BAO studies in eBOSS DR16 [10] and were referred as the London mocks. The methodology to produce this set of raw mocks is fully described in [40]. It starts with the generation of a correlated Gaussian multivariate overdensity field sampled in a three-dimensional mesh, given an input matter power spectrum, using the Cosmological Lofty Realizations package (CoLoRe [^6] https://github.com/damonge/CoLoRe[48]. The positions of the quasars in the mocks are assigned following an input number density $n(z)$ and bias $b(z)$, and placed in density peaks of the Gaussian random field above a given threshold via Poisson sampling. Density fields along the line of sight from the quasars to the center of the box, namely the skewers, are also extracted at this stage. The results are then post-processed with the LyaCoLoRe [^7] https://github.com/igmhub/LyaCoLoRe code, where small-scale fluctuations are added to the skewers in order to reproduce the variance of the Lyman-$\alpha$ forest in the data. A log-normal transformation is applied to the final Gaussian density field and a Fluctuating Gunn-Peterson approximation (FGPA) [11] is used to compute the optical depth on each simulated cell. Then, the effects of RSD are added by computing the radial velocity from the gradient of the Newtonian gravitational potential of the density field. This velocity is then used to shift the position of the optical depth in redshift space accordingly. Lastly, the optical depth $\tau$ is transformed into the transmitted flux by the equation $F=e^{-\tau}$. In addition, the LyaCoLoRe post-processing method includes the generation of transmission due to metals and also provides a catalog with the redshift and column densities of DLAs that are correlated with the density field, as briefly described in section 2.3.
@@ 2.2 Continuum templates
  • SIMQSO (default): Templates are generated using the simqso library [^9] https://github.com/imcgreer/simqso [52, 53], which contains a very broad set of tools to generate mock quasar spectra. The model used has two main components: a broken power-law continuum, and a set of Gaussian emission lines defined by their wavelength, equivalent width, and Gaussian RMS width ($\sigma$). To emulate the power-law continuum component, we randomly generate the slopes of the broken power-law continuum model following a Gaussian distribution centered at the slope value $m$ with dispersion $\sigma_{m}$ in a particular rest-frame wavelength region. We use the default values from simqso outside the Lyman-$\alpha$ region which are based in BOSS DR9: $m=-0.37$ at $5700\ \text{\AA}<\lambda<9730\ \text{\AA}$, $m=-1.70$ at $9730\ \text{\AA}<\lambda<22300\ \text{\AA}$ and -1.03 at $\lambda>22300\ \text{\AA}$, all with dispersion $\sigma_{m}=0.3$. For the Lyman-$\alpha$ forest region we tuned the slopes and dispersion to qualitatively reproduce the mean continuum and the dispersion (obtained from the first five components of a PCA analysis applied quasar spectra [54]) measured in the eBOSS DR16 quasar catalog. The resulting slopes are: $m=-1.5$ at $\lambda<1100\ \text{\AA}$ and $m=-0.5$ at $1100\ \text{\AA}<\lambda<5700\ \text{\AA}$, both with dispersion $\sigma_{m}=0.7$. To emulate the emission line component, we use the model described in detail in appendix A. The model combines emission lines within the Lyman-$\alpha$ forest region from the composite model of BOSS spectra [55] and emission lines outside the Lyman-$\alpha$ region from the simqso model. The emission line diversity is given mainly by the scatter in the equivalent width. Furthermore, the specific values for the mean equivalent width of some of the emission lines were adjusted so that the mean continuum resembles the results from DESI EDR+M2 data. Note that the emission line parameters used in this paper slightly differ from those used to generate the eBOSS DR16 mocks [10]. See the left panel of Figure 8 for a qualitative comparison of the mean continuum obtained in our analysis of the simulated spectra versus and the EDR+M2 data, which was also used to make the aforementioned adjustments of the slopes and the emission line parameters. Information about the continuum template generated for each mock quasar, such as the slopes and emission lines equivalent widths, are stored in one of the output "truth" files from quickquasars which are described in appendix C.
@@ 2.3 Astrophysical contaminants
Observed quasar spectra include features due to different astrophysical effects that contaminate the correlation function. In order to study their impact, these features may be added to the generated synthetic spectra. In the following we describe these contaminants.
2.3.1 Broad Absorption Lines systems
Quasars with broad ultraviolet (UV) absorption in their spectrum are called Broad Absorption Line (BAL) quasars. This absorption is produced by surrounding gas clouds near the quasar nucleus. BAL quasars are further classified on the basis of spectral lines that show troughs typically blueshifted from their rest-frame wavelength and with velocity widths larger than $2000\,\rm{km/s}$ [58]. For instance, HiBALs are BALs with absorption in high-ionization lines such as C IV(1549), LoBALs are those with absorption troughs from low-ionization features such as Mg II, whereas FeLoBALs exhibit absorption in lines such as Fe II [59, 60].
@@ 2.3.1 Broad Absorption Lines systems
2.3.2 High Columns Density systems
@@
  • Random. The simplest method is to add a random number of HCDs to the transmission at random wavelengths. In this case, the generated HCDs would not be correlated with the density field that was used to produce the raw mocks, which makes this method unsuitable for our purposes and therefore is not used for any of the mocks produced for this work. However, this method is being used in Ohio mocks to study the effect of DLAs in P1D analyses.
@@
2.3.3 Contamination from additional absorption transitions
Besides neutral-hydrogen, the IGM contains other absorbers, which we refer to as metals. Because our data analysis protocol transforms wavelengths to redshifts assuming Lyman-$\alpha$ absorption, neutral hydrogen and a metal occupying the same physical position are absorbed at different apparent redshifts, resulting in an apparent distance separation given by
@@ 2.3.3 Contamination from additional absorption transitions
where $D_{H}$ is the Hubble distance and $\Delta\lambda=\lambda_{m}-\lambda_{Ly\alpha}$ is the wavelength separation of the metal transition ($\lambda_{m}$) with respect to Lyman-$\alpha$. The auto and cross-correlation functions will then show peaks at $r_{\perp}\approx 0$ and the $r_{\parallel}$ scale of each metal of sufficiently strong absorption. The most important such metals are Si II(1260), Si III(1207), Si II(1193), and Si II(1190) corresponding respectively to $r_{\parallel}=104,\,-21,\,-54$ and $-61{\ \mathrm{Mpc/h}}$ at $z=2.35$.
In addition to this effect due to metal – neutral-hydrogen correlations, particularly strong absorbers can significantly affect the measured auto-correlation through their own (metal–metal) auto-correlation. The most important transitions are C IV(1548) and C IV(1550). However, these are not included in our mocks because absorption found within the Lyman-$\alpha$ forest region caused by these metals is produced by C IV at redshifts as low as $z\approx 1.32$, outside the range of the transmission files of the raw mocks.
@@
  • Added by quickquasars. In this case, we take the Lyman-$\alpha$ transmitted flux fraction at rest-frame as it is given in the transmission files. The optical depth of the Lyman-$\alpha$ absorption is then re-scaled by a factor $C_{m}$ according to the relative absorption strength of each metal with respect to Lyman-$\alpha$. The transmitted flux fraction of the metals is then computed as $F_{m}=e^{-C_{m}\tau_{Ly\alpha}}$. Finally, the absorption features are moved to their corresponding observed wavelengths according to the relation $\lambda_{obs,m}=(1+z_{abs})\lambda_{m}$, where $z_{abs}=\lambda/\lambda_{Ly\alpha}-1$ is the redshift at which the absorption occurs. In this method, the re-scaling of the optical depth is done after RSD effects have been applied to the Lyman-$\alpha$ absorption forcing the metals to have the same RSD parameter as Lyman-$\alpha$. The absorption amplitude of each metal transition strongly affects the value for its corresponding bias $b_{m}$ ($b_{i}$ defined in equation 2.1, for metals). In other words, the bias highly depends on the intensity of the optical depth of each independent transition and consequently on the value $C_{m}$ used in this method.
    For this simple simulation technique, $b_{m}$ and $C_{m}$ have a linear relation. This allows to find the $C_{m}$ needed to reproduce the measured biases of different eBOSS datasets. For this purpose, we made several realizations of the same mock including Lyman-$\alpha$ absorption, and absorption contamination due to Si II(1260), Si III(1207), Si II(1193), and Si II(1190). We varied the value of $C_{m}$ for each realization and measured the corresponding metal biases $b_{\eta,m}$. Then, we found their linear relation through a simple linear regression, and interpolated the $C_{m}$ value needed to obtain a target bias. In this case, we chose the absorption coefficient value that reproduced the results of eBOSS DR14 [8]. We performed this procedure on the Lyman-$\alpha$ transmission files of the LyaCoLoRe and Saclay mocks independently, the resulting $C_{m}$ values are shown in table 1. We set these values as default in quickquasars when using this method. However, other values of the $C_{m}$ parameters can be provided as an input to quickquasars if required. Note that the values for LyaCoLoRe are different that those of Saclay mocks in Table 1, this is expected since the values of the $C_{m}$ coefficients depend on the input details to construct the raw mocks, such as the flux-transmission distribution and the cell size used, to mention some.
    The $C_{m}$ values results were validated by computing the bias for a sample of ten LyaCoLoRe and ten Saclay mock realizations and were found to be statistically consistent with the target DR14 bias value of each metal. We refer the reader to [69] and [70] for further details about these results. The same values were also used to produce the mocks presented in the eBOSS DR16 analysis [10], also presenting consistency between mocks and observed data.
@@
    Si II(1190) $6.4239\times 10^{-4}$ $4.4960\times 10^{-4}$
    Si II(1193) $9.0776\times 10^{-4}$ $6.3540\times 10^{-4}$
    Si II(1260) $3.5420\times 10^{-4}$ $4.2504\times 10^{-4}$
    Si III(1207) $1.8919\times 10^{-3}$ $9.4595\times 10^{-4}$
  • From the transmission files. In this method the transmitted flux fraction of metals is computed during the generation of the Lyman-$\alpha$ transmission in the raw mocks. An example of this are the LyaCoLoRe raw mocks that included metal transmissions on their data products using the methodology explained in section 5.2 of [40], which we briefly describe here. Similarly to the method followed in quickquasars, here it is assumed that the optical depth of each additional metal is proportional to that of the Lyman-$\alpha$ absorption. Therefore for each metal we need to define a relative absorption strength $A_{m}$, so that $\tau_{m}=A_{m}\tau_{Ly\alpha}$ and a rest-frame wavelength at which the metal absorption will be included. Notice that in general, the absorption strength $A_{m}$ here is not the same as the coefficient $C_{m}$ mentioned above. However, the LyaCoLoRe mocks produced in [40] made use of the coefficients found using the quickquasars method, described above. [^11] All the relative absorption strength values used to produce the metal transmissions of the LyaCoLoRe mocks correspond to the $C_{m}$ coefficients found with the quickquasars method except for Si II(1190) where a value $A_{\text{Si~{}II(1190)}}=1.28478\times 10^{-4}$ was used to reach the target bias value of eBOSS DR12 $b_{\text{Si~{}II(1190)}}=-4.4\times 10^{-3}$ [6] (see Tables 2 and 3 in [40]). Finally, RSDs are applied and the transmitted flux fraction of each metal is computed separately. The aforementioned Si and Lyman-$\beta$ transmissions are then saved independently in the transmission files, which can then be read by quickquasars in the same way as the Lyman-$\alpha$ transmissions and added to the total transmitted flux.